# Decoding Open-Ended Information Seeking Goals from Eye Movements in Reading

### v2

@@ 5.2 Question Selection
We evaluate all the models on the question selection task and report selection accuracy. As described in Section˜4, the discriminative models assign a probability to a triplet $\langle T,q\in Q_{T},E_{P,T}\rangle$, and select the candidate question with the highest probability. For the generative models, we select the question with the highest model-assigned log-likelihood. In addition to an All overall three-way selection accuracy, we also report a breakdown of the model predictions into two fine-grained evaluations:
@@
Different critical spans This evaluation tests the model’s ability to distinguish between questions with different critical spans, reducing the selection problem to a binary decision between $q_{1,c_{1}}$ and $\{q_{2,c_{2}},q_{3,c_{2}}\}$. Note that random choice accuracy in this evaluation is approximately 55%, rather than 50%, due to a Monty Hall-type setup arising from grouping two questions under one critical span (see Table˜1).
@@
Same critical span This is a more challenging evaluation for differentiating between the two questions for $c_{2}$: $q_{2,c_{2}}$ versus $q_{3,c_{2}}$. Note that here we disregard the question whose critical span is $c_{1}$. The corresponding chance accuracy is 50%.
(a) All (b) Different critical spans (c) Same critical span
True
$\color[rgb]{0.9765625,0.56640625,0.40625}q_{1,c_{1}}$}
$\color[rgb]{0.08203125,0.49609375,0.49609375}q_{2,c_{2}}$}
$\color[rgb]{0.08203125,0.49609375,0.49609375}q_{3,c_{2}}$}
5.3 Question Reconstruction
@@
The generative models are further evaluated for the quality of the reconstructed questions. As with other generation tasks, evaluating reconstruction quality is a major challenge. Here, we use several heuristic evaluations for both the surface form and the semantic content of the generated questions.
@@
The question word We check if the generated question word matches the question word of the ground truth question. The possible question words are What, When, Where, Who, Whom, Which, Whose, Why, and How. An additional category is Other, reserved for all other words starting a question (7% of the questions). The agreement with the ground truth questions is measured using pairwise Cohen’s Kappa (Cohen, 1960).
@@
UIUC question category We test if the generated question has the same semantic category as the ground truth question, using the question-type taxonomy of Li & Roth (2002). We use GPT-4o (Hurst et al., 2024) to categorize the questions into the main taxonomy categories: entities, humans, numeric, and locations, and the subcategories of ‘descriptions’: manner, reasoning, definition, and description due to their high frequency in the data. We measure agreement in category assignments using pairwise Cohen’s Kappa. Section˜D.1 includes further details on the classification process and on the agreement of GPT-4o with manual human annotations.
@@
BertScore (Zhang et al., 2020) provides a cosine similarity score between the generated question and the true question based on contextual embeddings extracted from RoBERTa (Liu et al., 2019).
QA accuracy We introduce a downstream evaluation method with the following logic: the closer the generated question is to the true question, the easier it should be to select the correct answer for the true question when replaced with the generated question. We consider a question as a valid reconstruction of the true question if and only if, given this question, the text, and the four answers for the true question, the model selects the correct answer. We implement this evaluation using RoBERTa (Liu et al., 2019) fine-tuned for multiple choice QA on RACE (Lai et al., 2017), on the 86% of OneStopQA questions that RoBERTa answers correctly. We chose this model because OneStopQA is not in its training data, thus avoiding possible data contamination with newer models.
@@
To facilitate model performance interpretation, we introduce the following baseline questions:
Incorrect human questions The two additional human-composed questions for each text available in OneStop, one for a different critical span and one for the same critical span as the true question.
Arbitrary question from an LLM We generate an arbitrary question for the text with GPT-4o, using the prompt in Section˜D.2.
Question from a text-only question decoding model An additional question is generated by a Llama 3.1 model trained to decode the correct question without eye movement data.
@@ 6 Results
6.1 Question Selection
Table˜2 presents the aggregated test accuracy across the 10 data splits for the three-way ‘All’ selection, as well as a breakdown into questions with Different Spans, and questions with the Same Span evaluations. A division of the results into the New Text, New Participant, and New Text & Participant regimes is reported in Tables˜4, 5 and 6 in Appendix˜E. Validation set results are reported in Tables˜7, 8, 9 and 10 in Appendix˜E.
Table 2: Question selection test accuracy aggregated across 10 cross-validation splits, with 95% confidence intervals. A majority question choice is identical to chance because the data is balanced across questions. As the data is not i.i.d. (each participant reads multiple paragraphs, and each paragraph is read by multiple participants), we test for differences between models using a linear mixed effects model with random intercepts and slopes for participants and paragraphs: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$. The accuracy of all the models is compared to chance performance, where significant gains over this baseline are marked with ‘*’ $p<0.05$, ‘**’ $p<0.01$, and ‘***’ $p<0.001$. The best performing model in each evaluation regime is marked in bold. Significant drops compared to the best model are marked with ‘+’ $p<0.05$, ‘++’ $p<0.01$ and ‘+++’ $p<0.001$.
Model Type Model All
Different Spans
Same Span
Discriminative Chance / Majority Baseline $33.0\pm 0.4_{\text{+++}}$ $55.3\pm 0.4_{\text{+++}}$ $49.9\pm 0.4_{\text{+++}}$
@@ 6.1 Question Selection
Generative DalEye-LLaVA $33.5\pm 0.3_{\text{+++}}$ $57.0\pm 0.7^{**}_{\text{+++}}$ $49.4\pm 0.4_{\text{+++}}$
DalEye-Llama $39.2\pm 0.9^{***}_{\text{+++}}$ $61.4\pm 0.9^{***}_{\text{+++}}$ $52.5\pm 0.4^{***}_{\text{+++}}$
@@
Both Reading-Time Informed Embedding Similarity baseline models perform at chance level across all three evaluation regimes. A possible reason for this outcome lies in the inherent noise in the distribution of reading times of a single participant over a single paragraph, which deems heuristic methods that rely only on this distribution to be ineffective. Haller RNN and RoBERTEye-Fixations perform well above chance, and outperform the generative models on all the evaluations. In the generative models category, DalEye-Llama outperforms DalEye-LLaVA, where the latter is at chance level. Table˜11 in Section˜E.1.1 suggests that the alternative prompt and eye movements representations for DalEye-Llama weaken its performance. The advantage of the discriminative models over the generative models is not surprising, as differently from the discriminative models, the generative models are not explicitly trained on the question selection task.
@@
RoBERTEye-Fixations achieves the highest accuracy across all three evaluation regimes. In Table˜4 of Appendix˜E, we observe that this performance advantage is consistent across the New Text, New Participant, and New Text & Participant evaluations. Thus, the model is able to generalize not only to new participants but importantly also to new texts. Notably, it is also the only model that is well above chance in the Same Span regime with $57.3\%$ accuracy. On the one hand, this confirms that distinguishing between questions that ask about the same portion of the paragraph is indeed much more challenging than distinguishing between questions over different parts of the paragraph. At the same time, the above change performance of RoBERTEye-Fixations on the Same Span evaluation speaks to the fine-grained nature of the information that can be extracted about the reader’s information seeking goals from their eye movements alone.
6.2 Question Reconstruction
Based on the question selection results in Table˜2, we focus on the DalEye-Llama model in the evaluation of question reconstruction quality. Figure˜4 presents the evaluations in the New Participant and the New Text regimes. Table˜12 in Appendix˜E presents these results in numerical form, and additionally includes the New Text & Participant regime, which yields very similar results to the New Text regime. Figure˜7 in appendix˜E presents an example of DalEye-Llama outputs for the reconstruction task. Table˜13 in Appendix˜E presents reconstruction results for the alternative task prompt for DalEye-Llama, which yields similar results, and the alternative eye movement representations, which yield inferior results to those presented below.
@@ 6.2 Question Reconstruction
Figure 4: Question reconstruction evaluations of DalEye-Llama for (1) the identity of the generated question word, (2) UIUC semantic category of the question, (3) BLEU score, (4) BERTScore, (5) downstream QA accuracy based on the answer selection of a multiple-choice QA model. DalEye-Llama performance is benchmarked against two human-composed questions for a different and the same critical span, an arbitrary question generated with GPT-4o, and a question from a text-only Llama 3.1 model trained on the question decoding task using only the text. Presented are means with 95% confidence intervals (bootstrapped, $n=1000$).
Importantly, in the New Participant evaluation, DalEye-Llama generated questions outperform all the baseline questions on all five metrics. This suggests that DalEye-Llama generalizes relatively well to new participants when the text appeared in the training set. However, its performance drops both in absolute terms and relative to the baselines in the more challenging regime involving new texts. Notably, in this regime the incorrect human question - same critical span has the highest BLEU, BERTScore, and QA accuracy. This is likely the result of lexical overlap stemming from sharing the critical span with the true question. However, this span-sharing constraint may hinder this baseline in the Question word and Question category evaluations, as sharing a span tends to yield questions that differ in these regards. The human question over a different span, which tends to perform best in these evaluations, is not subject to this constraint.
@@ 7 Related Work
We build on two primary lines of work: (i) psycholinguistic investigation of goal driven reading; and (ii) predictive models of cognitive state from eye movements in reading.
7.1 Analyses of Goal Based Reading
@@
While most prior work in the psychology of reading and psycholinguistics focuses on ordinary reading for general comprehension Rayner et al. (2012), goal or task based reading has been acknowledged as a key frontier in this area (Radach & Kennedy, 2004). Several studies found differences in eye movements between ordinary reading and tasks that define a reading manner: skimming, speed reading and proofreading (Just et al., 1982; Kaakinen & Hyönä, 2010; Schotter et al., 2014; Strukelj & Niehorster, 2018; Chen et al., 2023). Additional studies have found differences between ordinary reading and tasks of target word search (Rayner & Raney, 1996; Rayner & Fischer, 1996), word verification (Radach et al., 2008), and topic search (White et al., 2015). Prior work also analyzed eye movements during human linguistic tasks, such as named entity annotation (Tomanek et al., 2010; Tokunaga et al., 2017). Differences in reading patterns were also found when readers adopt different perspectives on a given text (Kaakinen et al., 2002; Kaakinen & Hyönä, 2008), or given different learning goals (Rothkopf & Billington, 1979). Overall, these works consistently found a systematic influence of the reading task on the resulting reading patterns.
Our work builds most directly on Kaakinen et al. (2015), Malmaud et al. (2020), Hahn & Keller (2023), and Shubi & Berzak (2023), who compared eye movements in ordinary reading and open-ended information seeking. These studies showed task conditioning effects on reading patterns, especially as a function of the task-relevance of the textual information. Similarly to these studies, we formulate the information seeking task as a reading comprehension question specific to the text. We further leverage the previously observed task conditioning effects for automatic discrimination between different tasks over the same text.
7.2 Decoding Cognitive State from Eye Movements
@@
Multimodal modeling of eye movements and text has been a growing area of research in recent years (Reich et al., 2025), with applications ranging from improving NLP (e.g. Klerke et al., 2016; Mishra et al., 2016; Barrett & Hollenstein, 2020; Sood et al., 2020; Deng et al., 2024; López-Cardona et al., 2025) to prediction of eye movement trajectories (e.g. Hollenstein et al., 2021a; Bolliger et al., 2023; Deng et al., 2023). Within this broad line of research, several studies focused on cognitive state decoding: prediction of language-related aspects of the reader and their cognitive state from their eye movements while reading. These include studies on prediction of the reader’s linguistic background (Berzak et al., 2017; Reich et al., 2022; Skerath et al., 2023), language proficiency (Berzak et al., 2018), reading comprehension (Ahn et al., 2020; Reich et al., 2022; Shubi et al., 2024), and subjective readability (Reich et al., 2022).
@@
Most notably, two studies addressed decoding of the reader’s goals. Hollenstein et al. (2021b) classified ordinary reading for comprehension versus annotation of semantic relations in single sentences using the ZuCo corpus (Hollenstein et al., 2020; 2023). Shubi et al. (2025) classified ordinary reading versus information seeking using the OneStop dataset. Importantly, in both studies, the decoding tasks are procedural, where the goal is to distinguish between a small set of pre-defined manners of reading that are not specific to a particular text. In contrast, our study addresses a related but conceptually and practically different task, which is not procedural, but rather semantic, decoding reading tasks specific to the text, within the information seeking regime. Instead of a small number of categories that can apply to any text, we have hundreds of text-specific tasks that can be arbitrary in nature. See further discussion on the procedural versus semantic tasks distinction in Shubi & Berzak (2023). Furthermore, these and other studies on cognitive state decoding typically use discriminative, encoder-based models that are limited to classification tasks. Here, we also integrate eye movements into generative, decoder-based language models, and investigate both textual- and embedding-based methods for conditioning text generation on eye movements in reading.

### v3

@@ 5.2 Goal Selection
As described in Section˜4, the discriminative models assign a probability to a triplet $\langle T,E_{P,T},q\in Q_{T}\rangle$, and select the candidate question with the highest probability. We report a three-way selection accuracy All, as well as accuracy for two specific cases of interest:
@@
Different critical spans This evaluation tests the ability to distinguish between questions with disjoint critical spans. We pose the selection problem as a binary decision between $q_{1,c_{1}}$ and $\{q_{2,c_{2}},q_{3,c_{2}}\}$. Random choice accuracy in this evaluation is approximately 55%, rather than 50%, due to a Monty Hall-type setup arising from grouping two questions under one critical span (see Appendix˜E).
@@
Same critical span This is a more challenging evaluation for differentiating between the two questions with the same critical span $c_{2}$: $q_{2,c_{2}}$ versus $q_{3,c_{2}}$. Note that here we disregard the question whose critical span is $c_{1}$. The corresponding chance accuracy is 50%.
5.3 Goal Reconstruction
@@
The generative models are evaluated on the quality of the reconstructed questions. As with other generation tasks, evaluating reconstruction quality is a major challenge. Here, we use several heuristic evaluations for both the surface form and the semantic content of the generated questions.
@@
Question word accuracy We test whether the generated question word matches the question word of the ground truth question. The possible question words are What, When, Where, Who, Whom, Which, Whose, Why, and How. An additional category is Other, reserved for all other words starting a question (7% of the questions).
@@
UIUC question category accuracy We check whether the generated question has the same semantic category as the ground truth question according to the question-type taxonomy of Li and Roth (2002). We use GPT-4o (Hurst et al., 2024) to categorize the questions into the main taxonomy categories: entity, human, numeric, and location, and the subcategories of ‘description’: manner, reasoning, definition, and description. Section˜F.1 includes further details on the classification process and on the agreement of GPT-4o with manual human annotations.
@@
BERTScore (Zhang et al., 2020) A cosine similarity score between the generated question and the true question based on contextual embeddings extracted from RoBERTa.
QA accuracy A downstream evaluation with the following logic: the closer the generated question is to the true question, the easier it should be to choose the correct answer for the true question given the reconstructed question. We consider a question as a valid reconstruction of the true question if and only if, given this question, the text, and the four answers for the true question, the model selects the correct answer. We implement this evaluation using RoBERTa fine-tuned for multiple choice QA on RACE (Lai et al., 2017), on the 86% of OneStopQA questions that RoBERTa answers correctly.
@@
To facilitate model performance interpretation, we introduce the following baseline questions which are composed without eye movement information.
Human incorrect questions The two additional human-composed questions for each text, one with a different and one with the same critical span as the true question (Section˜2). These questions are guaranteed to fulfill two key criteria: they are well-formed and can be answered from the passage.
LLM arbitrary questions An arbitrary question for each passage generated with the most recent frontier LLM, Gemini-3-Pro-Preview, using the prompt presented in Section˜F.2.
LLM questions from text-only question decoding models Questions from Llama 3.1 and GPT-4o-mini finetuned to decode the question in using the same experimental setup as DalEye-Llama and DalEye-GPT, but without eye movements data.
The questions generated from the LLM baselines enable quantifying the contribution of eye movement information in each of our generative models, above and beyond textual heuristics. This includes effects of possible model exposure to the textual materials of OneStopQA during pretraining.
@@ 6 Results
6.1 Goal Selection
Table 1: Goal selection test accuracy aggregated over 10 data splits, with 95% confidence intervals. A majority choice is identical to chance because the data is balanced across questions. As the data is not i.i.d. (each participant reads multiple paragraphs, and each paragraph is read by multiple participants), we test for differences between models using a linear mixed effects model with random intercepts and slopes for participants and paragraphs: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$. Significant gains over chance performance are marked with ‘*’ $p<0.05$, ‘**’ $p<0.01$, and ‘***’ $p<0.001$. The best performing model in each evaluation is marked in bold. Significant drops compared to the best model are marked with ‘+++’ $p<0.001$.
Model All Different Spans Same Span
Chance / Majority Baseline $33.0\pm 0.4_{\text{+++}}$ $55.3\pm 0.4_{\text{+++}}$ $49.9\pm 0.4_{\text{+++}}$
@@ 6.1 Goal Selection
Table˜1 presents the aggregated test accuracy across the 10 data splits for the three-way All goal selection, and a breakdown into questions with Different Spans, and the Same Span evaluations. A division of the results into the New Participant, New Text, and New Text & Participant regimes is reported in Tables˜5, 6 and 7 in Section˜G.1.1. Validation set results are reported in Tables˜8, 9, 10 and 11 in Section˜G.1.1. A secondary evaluation of the open generative models DalEye-Llama and DalEye-LLaVA on the question selection task is reported in Section˜G.1.2.
@@
Both Reading-Time Informed Embedding Similarity baseline models perform at chance level across all three evaluation regimes, with the exception of RT Similarity to Question–Word Similarities in the All condition, where performance is marginally above chance. A possible reason for this outcome lies in the inherent noise in the distribution of reading times of a single participant over a single paragraph, which deems heuristic methods that rely only on this distribution to be ineffective. Haller RNN and RoBERTEye-Fixations on the other hand perform well above chance.
@@
RoBERTEye-Fixations achieves the highest accuracy across all three evaluation regimes. In Table˜5 of Section˜G.1.1, we observe that this performance advantage is consistent across the New Participant, New Text, and New Text & Participant evaluations. Thus, the model is able to generalize not only to new participants but importantly also to new texts. Notably, in addition to achieving $70.9\%$ accuracy on the Different Spans evaluation, it is also the only model that is well above chance in the Same Span regime with $57.3\%$ accuracy. While these results confirm that distinguishing between questions that ask about the same portion of the text is indeed much more challenging than distinguishing between questions over different parts of the text, the above-chance performance of RoBERTEye-Fixations on the Same Span evaluation speaks to the fine-grained nature of the information that can be extracted about the reader’s information seeking goals from their eye movements alone. Section˜G.1.3 reports a feature ablation analysis of RoBERTEye-Fixations, which suggests that the model relies much more heavily on the ordering of word embeddings according to the sequence of fixations than on any additional features of each individual fixation.
Linking Model Performance to Cognitive Indicators of Information Seeking
To tie automated goal decoding to cognitive aspects of gaze behavior during information seeking, we present a model performance analysis of RoBERTEye-Fixations, the top performing model in the goal selection task, based on the framework proposed by Shubi et al. (2025). The analysis examines how features of the experimental trial that are pertinent to cognitive theory of information seeking (e.g. Kaakinen et al. (2002); Hahn and Keller (2023); Shubi and Berzak (2023)) and are not provided to the model, affect model performance. Specifically, it uses the following linear mixed effects model which accounts for the non i.i.d structure of the data: $P(\text{is\_correct})\sim\text{feat}_{1}+\cdots+\text{feat}_{n}+(1\mid\text{participant})+(1\mid\text{paragraph})$, where PP(is_correct) is the probability assigned by RoBERTEye-Fixations to the correct question given an input of a paragraph and eye movements. We use the 10 trial features from Shubi et al. (2025) which capture aspects of the participant’s interaction with the textual item, and the item itself, and further include the lexical overlap between the question and critical span, quantified with ROUGE-1-precision. Figure˜4 depicts the resulting model coefficients.
Figure 4: Coefficients from a mixed-effects model which predicts the probability assigned by RoBERTEye-Fixations to the correct question from 11 features of the experimental trial. Predictors are z-normalized, to make the coefficient magnitudes comparable. Coefficient statistical significance after a 11x Bonferroni correction is marked with ‘*’ $p<0.05$, ‘**’ $p<0.01$, and ‘***’ $p<0.001$.
Participant features over the textual item Prior studies in psycholinguistics have observed shorter reading times on goal irrelevant compared to goal relevant information, which was linked to rational processing during information seeking (Kaakinen et al., 2002; Hahn and Keller, 2023; Shubi and Berzak, 2023). Intriguingly, we find that the extent to which this holds affects model performance in a highly interpretable manner: longer reading times (per word Total Fixation Duration) before and after the critical span hurt model performance ($p<10^{-63}$ and $p<10^{-71}$ respectively), while longer reading times within the critical span improve it ($p<10^{-275}$). In other words, the more goal directed and resource efficient the reader is, the easier it is to identify their goal automatically. The paragraph’s position in the experiment (1-54) is also positively correlated with model performance ($p<10^{-3}$), likely due to participants’ adaptation to the task over time (e.g. Wells et al., 2009; Chromỳ and Tomaschek, 2024). Better reading comprehension is similarly positively associated with model performance ($p<0.01$), in line with work linking better time division between goal relevant and irrelevant information and reading comprehension performance (Shubi and Berzak, 2023).
Textual item features Longer paragraphs are associated with improved model performance ($p<10^{-6}$), possibly as they allow better separation between question relevant and question irrelevant material, as well as between critical spans of different questions. Longer critical spans, which potentially support more robustness in the extraction of critical span related eye movements, are also positively correlated with model performance ($p<10^{-8}$). Interestingly, this result complements an opposite trend for models that are trained to distinguish between information seeking and reading for comprehension without receiving the question ahead of time (Shubi et al., 2025), where longer critical spans make it harder for models to distinguish these two reading regimes. Similarly to the results in Shubi et al. (2025), paragraph difficulty level (original or simplified), and critical span location do not have a significant effect on the model’s performance. Finally, we do not find evidence for an influence of the lexical overlap of the critical span with the question. Overall, these findings provide converging evidence and expand prior observations from the psycholinguistic literature on information seeking using the developed model as an analytical tool.
6.2 Goal Reconstruction
@@
Figure 5: Goal reconstruction from eye movements evaluations of the finetuned models DalEye-Llama and DalEye-GPT, as well as zero-shot and few-shot Gemini for (1) the identity of the generated question word, (2) UIUC semantic category of the question, (3) BLEU score, (4) BERTScore, (5) downstream QA accuracy based on the answer selection of a multiple-choice QA model. The models are benchmarked against five baselines that do not include eye movements. Two baselines are human-composed questions for a different and the same critical span. Three additional baselines are LLM-generated questions: an arbitrary question about the text generated with Gemini 3, and questions from text-only variants of DalEye-Llama and DalEye-GPT finetuned on the question decoding task using only the text. Results are aggregated over 10 data splits. Presented are intercept coefficients with 95% confidence intervals from linear mixed effects models with random intercepts for participants and paragraphs: $measure\sim 1+(1\mid participant)+(1\mid paragraph)$ for eye movement models and $measure\sim 1+(1\mid paragraph)$ for the baselines. The highest score in each combination of evaluation measure and evaluation regime is marked with $\star$.
Figure˜5 presents the goal reconstruction evaluations in the New Participant and the New Text evaluation regimes. Tables˜14 and 15 in Section˜G.2 present these results numerically, and also include the New Text & Participant regime, which yields similar results to the New Text regime. In the New Participant evaluation, DalEye-Llama, DalEye-GPT and Gemini few-shot generated questions outperform all the baseline questions on all five evaluation measures ($p<0.001$ in all cases) [^2] Performance differences are tested using: $measure\sim model+(1\mid participant)+(1\mid paragraph)$., except for Gemini few-shot in the Question Category evaluation. These performance gains suggest that these models generalize relatively well to new participants when the text and possible questions appeared in the training set. Gemini few-shot is the top performer in this regime, with the exception of Question Category where DalEye-Llama performs best. Notably, the text-only version of DalEye-Llama, Llama 3.1 finetuned for question reconstruction without eye movements, is the strongest baseline for this evaluation regime across all five evaluation measures.
The performance of DalEye-Llama, DalEye-GPT, and Gemini few-shot drops considerably in the New Text regime, in absolute terms and relative to the baselines, suggesting weaker generalization to new texts and questions. However, even in this challenging evaluation regime, all three models consistently outperform their corresponding text-only or arbitrary counterparts, except for DalEye-Llama on the Question Word and Question Category evaluation. This indicates that eye movements are valuable for the generation task even when the textual item is novel. The best performing model differs by evaluation measure: DalEye-Llama on Question Word, DalEye-GPT on Question Category, and Gemini few-shot on QA accuracy. The Incorrect Human (same critical span) questions achieve the best results on BLEU and BERTScore, likely due to lexical overlap and semantic proximity to the correct question, stemming from sharing the same critical span. The performance of Gemini zero-shot is similar across both regimes, because similarly to the human baselines, this model is not finetuned, and it tends to perform poorly in both regimes.
Table˜16 in Section˜G.2 presents examples of questions generated by DalEye-Llama, DalEye-GPT, and Gemini. Table˜18 in Section˜G.2 presents experimental results for DalEye-Llama using a differently worded task prompt, which yields similar results, as well as two alternative textual representations of the eye movement sequence, one based on word-level features and another combining word- and fixation-level features. The alternative representations yield inferior results to those presented above.
@@ 7 Related Work
Analyses of Goal Based Reading While most prior work in psycholinguistics focuses on reading for comprehension (Rayner et al., 2012), goal or task based reading has been acknowledged as a key frontier in this area (Radach and Kennedy, 2004). Several studies found differences in eye movements between reading for comprehension and tasks that define a reading manner: skimming, speed reading and proofreading (Just et al., 1982; Kaakinen and Hyönä, 2010; Schotter et al., 2014; Strukelj and Niehorster, 2018; Chen et al., 2023). Additional studies found differences between reading for comprehension and tasks of target word search (Rayner and Raney, 1996; Rayner and Fischer, 1996), word verification (Radach et al., 2008), and topic search (White et al., 2015). Prior work also analyzed eye movements during linguistic annotation tasks (Tomanek et al., 2010; Tokunaga et al., 2017). Differences in reading patterns were also found when readers adopt different perspectives on a text (Kaakinen et al., 2002; Kaakinen and Hyönä, 2008), or given different learning goals (Rothkopf and Billington, 1979). Overall, these works consistently found a systematic influence of the reading task on reading patterns. We build most directly on Kaakinen et al. (2015), Malmaud et al. (2020), Hahn and Keller (2023), and Shubi and Berzak (2023), who compared eye movements in reading for comprehension and open-ended information seeking. These studies showed task conditioning effects on reading patterns, especially as a function of the task-relevance of the textual information. Similarly to these studies, we formulate the information seeking task as a reading comprehension question.
@@
Decoding Cognitive State from Eye Movements Multimodal modeling of eye movements and text has been a growing research area in recent years (Reich et al., 2025), with applications ranging from improving NLP (e.g. Klerke et al., 2016; Mishra et al., 2016; Barrett and Hollenstein, 2020; Sood et al., 2020; Khurana et al., 2023; Deng et al., 2024; López-Cardona et al., 2025) to prediction of eye movements (e.g. Hollenstein et al., 2021a; Bolliger et al., 2023; Deng et al., 2023; Chen et al., 2024). Within this line of research, several studies focused on cognitive state decoding: prediction of language-related aspects of the reader and their cognitive state from eye movements. These include studies on prediction of the reader’s linguistic background (Berzak et al., 2017; Reich et al., 2022; Skerath et al., 2023), language proficiency (Berzak et al., 2018), reading comprehension (Ahn et al., 2020; Reich et al., 2022; Shubi et al., 2024), subjective readability (Reich et al., 2022) and perceived relevance (Bhattacharya et al., 2020b; a).
@@
Decoding Reading Goals Most notably, two studies addressed decoding of the reader’s goals. Hollenstein et al. (2021b) classified reading for comprehension versus annotation of semantic relations, Shubi et al. (2025) classified reading for comprehension versus information seeking. Importantly, in both studies, the decoding tasks are procedural, where the goal is to distinguish between a small set of pre-defined manners of reading that are not specific to a particular text. In contrast, our study addresses a related but conceptually and practically different task, which is semantic; instead of a small number of categories that can apply to any text, we have hundreds of text-specific goals that can be arbitrary in nature. See further discussion on the procedural versus semantic distinction in Shubi and Berzak (2023). Furthermore, these and other studies on cognitive state decoding typically use discriminative, encoder-based models that are limited to classification tasks. Here, we also develop generative, decoder-based language models. We envision that such models will have broad relevance for future eye movement conditioned text generation tasks.